# Extended mean field control problems with constraints: The generalized Fritz-John conditions and Lagrangian method

### v1

@@ 2.2 The generalized FJ optimality condition on Banach spaces
In this subsection, we establish a necessary optimality condition for problem (1) with Banach space valued constraints, namely the generalized FJ optimality condition, which is customized for the application to the C-(E)MFC problem in Section 3.
@@
Before proving Theorem 2.7, let us first recall the Ekeland’s variation principle as below.
Lemma 2.9 (Ekeland’s variation principle [23]).
Let $(V,d)$ be a complete metric space and $F:C\subset V\mapsto\mathbb{R}$ be a lower semi-continuous function that is bounded from below. Assume that there exist some $\widehat{x}\in C$ and $\epsilon>0$ such that $F(\widehat{x})\leq\inf_{v\in C}F(v)+\epsilon$. Then, there exists a $z_{\epsilon}\in C$ such that (i) $F(z_{\epsilon})\leq F(\widehat{x})$, (ii) $d(\widehat{x},z_{\epsilon})\leq\sqrt{\epsilon}$, (iii) $F(v)+\sqrt{\epsilon}d(v,z_{\epsilon})\geq F(z_{\epsilon})$ for all $v\in C$.
@@
Let $\epsilon\in(0,1)$ and recall $F:\mathtt{X}\to\mathbb{R}$ given in Lemma 2.9 that, for $x\in C$,
@@
Then, $x\mapsto F(x)$ is locally Lipschitz and bounded below by 00. As $\widehat{x}$ solves the problem (1) locally, we have
$F(\widehat{x})$ $=\max\left\{\epsilon,~{}\!\!\!\!\sup_{\begin{subarray}{c}y_{i}^{*}\in K_{i}^{+}\\ \lVert y_{i}^{*}\rVert_{\mathtt{Y}_{i}^{*}}=1\end{subarray}}\langle g_{i}(\widehat{x}),y_{i}^{*}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}},~{}0;~{}i\in I,~{}j\in J\right\}=\epsilon,$
and $\widehat{x}$ is an $\epsilon$-minimum of $x\mapsto F(x)$ over $C$. Thus, we can apply Lemma 2.9 to conclude the existence of $z_{\epsilon}\in C$ such that the assertions in Lemma 2.9 hold.
@@
intersects $\sqrt{\epsilon}B_{1}^{*}$ where $B_{1}^{*}$ is the unit ball in $\mathtt{X}^{*}$, and $\phi_{i}^{\epsilon}=0$ if $g_{i}(z_{\epsilon})\leq_{K_{i}}0$ for $i\in I$ and $\psi_{j}^{\epsilon}=0$ if $h_{j}(z_{\epsilon})=0$ for $j\in J$. It is easy to verify that $x\mapsto\lVert h_{j}(x)\rVert_{\mathtt{H}_{j}}$ is continuously Fréchet differentiable and $D\lVert h_{j}(x)\rVert_{\mathtt{H}_{j}}=(h_{j}(x)/\lVert h_{j}(x)\rVert_{\mathtt{H}_{j}})\circ Dh_{j}(x)$ when $h_{j}(x)\neq 0$ for $i\in I$. Thus, let us set
$\mu_{j}^{\epsilon}:=-(h_{j}(z_{\epsilon})/\lVert h_{j}(z_{\epsilon})\rVert_{\mathtt{H}_{j}})\psi_{j}^{\epsilon}\in\mathtt{H}_{j}$
when $h_{j}(z_{\epsilon})=0$, and set $\mu_{j}^{\epsilon}:=0$ otherwise, for $j\in J$.
On the other hand, we have
$D\langle g_{i}(x),y_{i}^{*}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}=y_{i}^{*}\circ Dg_{i}(x)\in\mathtt{X}^{*}$
for $x\in\mathtt{X}$. We claim that the set $\{y_{i}^{*}\circ Dg_{i}(x);~{}y_{i}^{*}\in M(x)\}$ is closed and convex, where $M_{i}(x)$ is the maximizer set given by
@@
As $\langle g_{i}(x),y_{i}^{*}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}$ is weak-* continuous in $y_{i}^{*}$ for any $x\in\mathtt{X}$ and $\{y_{i}^{*}\in K_{i}^{+};~{}\lVert y_{i}^{*}\rVert_{\mathtt{Y}_{i}^{*}}=1\}$ is weak-* compact due to Alaogu’s theorem, $M_{i}(x)$ is nonempty, and moreover it can be easily verified to be convex and weak-* closed (and hence closed). The convexity of $\{y_{i}^{*}\circ Dg_{i}(x);~{}y_{i}^{*}\in M_{i}(x)\}$ follows from $Dg_{i}(x)\in\mathcal{L}(\mathtt{X},\mathtt{Y}_{i})$ and the convexity of $M_{i}(x)$. To show the closedness, let us assume that $y_{i}^{n^{*}}\circ Dg_{i}(x)\to x^{*}$ in $\mathtt{X}^{*}$ with $y_{i}^{n^{*}}\in M_{i}(x)$. As the latter set is bounded, according to Lemma 2.5-(i), there exists a subsequence $(y_{i}^{n_{k}^{*}})$ of $(y_{i}^{n^{*}})$ , weakly- converging to some $y^{*}\in\mathtt{Y}_{i}^{*}$ (in fact $y^{*}\in M_{i}(x)$ in view that $M_{i}(x)$ is weak- closed). Thus, we can conclude that
@@ 2.3 Constraint Qualifications
Next, we also provide a sufficient condition for the KKT optimality condition to hold. Let $\Pi_{1}(C):=\{\Pi_{1}(x);~{}x\in C\}$. Then
@@
Proof.
In above, the equality can only be attained when $\lambda_{i}=0$ for $i\in I$. Thus, the equality (7) can be translated into
$\sum_{j\in J}\mu_{j}\circ Dh_{j}(\widehat{x})=\xi\in N_{C}(\widehat{x}).$
Here, $\xi=(\xi_{1},\xi_{2})\in\mathtt{X}_{1}^{*}\times\mathtt{X}_{2}^{*}$ as the dual space of a product space is the product of their dual spaces. Hence, we have $\sum_{j\in J}\mu_{j}\circ D_{x_{1}}h_{j}(\widehat{x})=\xi_{1}$. Moreover, by the definition of normal cone, we can conclude that by choosing $c=(c_{1},\widehat{x}_{2})\in C\subset\mathtt{X}_{1}\times\mathtt{X}_{2}$, where we recall $\widehat{x}=(\widehat{x}_{1},\widehat{x}_{2})\in C\subset\mathtt{X}$, we have
$\sum_{j\in J}\mu_{j}\circ D_{x_{1}}h_{j}(\widehat{x}_{1})=\xi_{1}=0.$ (9)
We argue by contradiction to show that $\lambda_{i}=0$ for $i\in I$. To this end, we first suppose that $\ker(\lambda_{1})^{c}\neq\varnothing$. Then, by the surjection introduced in LICQ($\widehat{x}$), there exists an $x_{0}\in\mathtt{X}_{1}$ such that $(D_{x_{1}}g_{i}(\widehat{x}_{1})(x_{0}),D_{x_{1}}h_{j}(\widehat{x}_{1})(x_{1}))_{i\in I,j\in J}$ belongs to the nonempty set $\ker(\lambda_{1})^{c}\times\prod_{i\in I\backslash\{1\}}\ker(\lambda_{i})\times\prod_{j\in J}\ker(\mu_{j})$, which yields that
@@ 3 Extended MFC Problems with Constraints
In this section, we apply the newly obtained FJ optimality condition in Theorem 2.7 to establish the SMP for some C-(E)MFC problems in which we can encompass both dynamic expectations constraints and mixed state-control constraints. To this end, we first formulate our C-(E)MFC problem and introduce some basic spaces and notations as preparations.
Let $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a filtered probability space with the filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\in[0,T]}$ satisfying the usual conditions. For $r,l\in\mathbb{N}$, let $(W_{t})_{t\in[0,T]}$ be a standard $r$-dimensional $(\mathbb{P},\mathbb{F})$-Brownian motion and denote $U\subset\mathbb{R}^{l}$ as the closed and convex control space.
We define $\mathcal{X}$ as the space $L^{2}([0,T];\mathbb{R}^{l})/\mathord{\sim}$ equipped with the inner product $\langle\gamma^{1},\gamma^{2}\rangle_{\mathcal{X}}=\int_{0}^{T}\langle\gamma_{t}^{1},\gamma_{t}^{2}\rangle\mathrm{d}t$. Then, $(\mathcal{X},\langle\cdot,\cdot\rangle_{\mathcal{X}})$ is a Hilbert space. For any $\gamma=(\gamma_{t})_{t\in[0,T]}\in\mathcal{X}$, we introduce, for fixed $t\in[0,T]$,
@@
where $\beta^{1}=\beta^{2}$ means that $\beta^{1}(\omega)=\beta^{2}(\omega)$ holds in $\mathcal{X}$ except for a $\mathbb{P}$-null set. It can be easily verified that $\mathscr{U}[0,T]$ is closed and convex in $(\mathbb{L}_{\mathbb{F}}^{2},\langle\cdot,\cdot\rangle_{\mathbb{L}_{\mathbb{F}}^{2}})$. We can observe that $\beta\in L^{2}(\Omega;\mathcal{X})$ for all $\beta\in\mathbb{L}_{\mathbb{F}}^{2}$, and moreover, we have
$\|\beta\|_{\mathbb{L}_{\mathbb{F}}^{2}}^{2}:=\mathbb{E}\left[\|\beta\|_{\mathcal{X}}^{2}\right]=\mathbb{E}\left[\langle\beta,\beta\rangle_{\mathcal{X}}\right].$ (12)
Let us define $\mathbb{M}^{2}$ ($\subset\mathbb{L}_{\mathbb{F}}^{2}$) as the set of $\mathbb{F}$-adapted (continuous) $\mathbb{R}^{n}$-valued process $\eta=(\eta_{t})_{t\in[0,T]}$ satisfying the Itô’s representation that
@@ 3.1 Partial L-derivative
Let $w:\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})\mapsto\mathbb{R}$. We use $\rho$ to represent a generic element in $\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})$, while $\mu\in\mathcal{P}_{2}(\mathbb{R}^{n})$ and $\nu\in\mathcal{P}_{2}(\mathbb{R}^{l})$ denote its respective marginals. For a probability space $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{P}})$, define the lift $\tilde{w}:L^{2}((\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{P}});\mathbb{R}^{n}\times\mathbb{R}^{l})$ of the element $w$ by
$\tilde{w}(\tilde{X},\tilde{\alpha})=w\left(\mathcal{L}(\tilde{X},\tilde{\alpha})\right),$ (16)
where $\mathcal{L}(\tilde{X},\tilde{\alpha})$ denotes the joint distribution of $(\tilde{X},\tilde{\alpha})$ under $\tilde{P}$. Then, the L-derivative is defined by
@@
3.2 Formulation of C-(E)MFC problems
In this section, we formulate our C-(E)MFC problems. Let $(b,\sigma):[0,T]\times\mathbb{R}^{n}\times{\cal P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})\times\mathbb{R}^{l}\mapsto\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ be measurable mappings. For any $\alpha=(\alpha_{t})_{t\in[0,T]}\in\mathscr{U}[0,T]$, consider the following controlled Mckean-Vlasov SDE that
$\mathrm{d}X_{t}^{\alpha}$ $=b(t,X_{t}^{\alpha},\mathcal{L}(X_{t}^{\alpha},\alpha_{t}),\alpha_{t})\mathrm{d}t+\sigma(t,X_{t}^{\alpha},\mathcal{L}(X_{t}^{\alpha},\alpha_{t}),\alpha_{t})\mathrm{d}W_{t},$
$X_{0}^{\alpha}$ $=\kappa\in L^{2}((\Omega,\mathcal{F}_{0},\mathbb{P});\mathbb{R}^{n}),$ (17)
where $\mathcal{L}(X_{t}^{\alpha},\alpha_{t})$ denotes the joint distribution of $(X_{t}^{\alpha},\alpha_{t})$ for $t\in[0,T]$.
Let integer $1\leq m_{1}\leq m$. For $\alpha\in\mathscr{U}[0,T]$ and $\phi^{i}:[0,T]\times\mathbb{R}^{n}\times\mathcal{X}\mapsto\mathbb{R}$ with $i\in I_{1}:=\{1,2,\ldots,m_{1}\}$, we introduce the dynamic expectation constraints in the sense that, for $i\in I_{1}$,
$\mathbb{E}\left[\phi^{i}\left(t,X_{t}^{\alpha},\widetilde{\alpha}^{t}\right)\right]\geq 0,\quad\forall t\in[0,T].$ (18)
For $\psi^{j}:[0,T]\times C([0,T];\mathbb{R}^{n})\times\mathbb{R}^{l}\mapsto\mathbb{R}$ with $j\in I_{2}:=I\setminus I_{1}$, the dynamic state-control constraint is formulated by, for $j\in I_{2}$,
$\psi^{j}\left(t,X_{t}^{\alpha},\alpha_{t}\right)\geq 0,\quad\mathrm{d}t\times\mathrm{d}\mathbb{P}\text{-}\mathrm{a.s.}.$ (19)
Here, $X^{\alpha}=(X_{t}^{\alpha})_{t\in[0,T]}$ obeys the controlled Mckean-Vlasov SDE (3.2) and the process $\widetilde{\alpha}^{t}$ is defined by (10). Then, we introduce the admissible control set $\mathscr{U}_{ad}[0,T]$ as the set of processes $\alpha\in\mathscr{U}[0,T]$ such that there exists $X^{\alpha}=(X_{t}^{\alpha})_{t\in[0,T]}\in\mathbb{M}^{2}$ satisfying (3.2), (18) and (19). Based on the preparations provided above, we introduce the cost functional as follows, for $\alpha=(\alpha_{t})_{t\in[0,T]}\in\mathscr{U}_{ad}[0,T]\neq\varnothing$,

### v2

@@ 2.2 The generalized FJ optimality condition on Banach spaces
In this subsection, we establish a necessary optimality condition for problem (1) with Banach space valued constraints, namely the generalized FJ optimality condition.
@@
Let $\epsilon\in(0,1)$ and $F:\mathtt{X}\to\mathbb{R}$ be defined by, for $x\in C$,
@@
Then, $x\mapsto F(x)$ is locally Lipschitz and bounded below by 00. As $\widehat{x}$ solves the problem (1) locally, we have $F(\widehat{x})=\epsilon$, and $\widehat{x}$ is an $\epsilon$-minimum of $x\mapsto F(x)$ over $C$. Thus, we can apply Ekeland’s variation principle given by Theorem 1.1 in [23] to conclude the existence of $z_{\epsilon}\in C$ such that the corresponding assertions in Ekeland’s variation principle hold.
@@
intersects $\sqrt{\epsilon}B_{1}^{*}$ where $B_{1}^{*}$ is the unit ball in $\mathtt{X}^{*}$, and $\phi_{i}^{\epsilon}=0$ if $g_{i}(z_{\epsilon})\leq_{K_{i}}0$ for $i\in I$ and $\psi_{j}^{\epsilon}=0$ if $h_{j}(z_{\epsilon})=0$ for $j\in J$. It is easy to verify that $x\mapsto\lVert h_{j}(x)\rVert_{\mathtt{H}_{j}}$ is continuously Fréchet differentiable and $D\lVert h_{j}(x)\rVert_{\mathtt{H}_{j}}=(h_{j}(x)/\lVert h_{j}(x)\rVert_{\mathtt{H}_{j}})\circ Dh_{j}(x)$ when $h_{j}(x)\neq 0$ for $i\in I$. Thus, let us set $\mu_{j}^{\epsilon}:=-(h_{j}(z_{\epsilon})/\lVert h_{j}(z_{\epsilon})\rVert_{\mathtt{H}_{j}})\psi_{j}^{\epsilon}\in\mathtt{H}_{j}$ when $h_{j}(z_{\epsilon})=0$, and set $\mu_{j}^{\epsilon}:=0$ otherwise, for $j\in J$. On the other hand, we have $D\langle g_{i}(x),y_{i}^{*}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}=y_{i}^{*}\circ Dg_{i}(x)\in\mathtt{X}^{*}$ for $x\in\mathtt{X}$. We claim that the set $\{y_{i}^{*}\circ Dg_{i}(x);~{}y_{i}^{*}\in M(x)\}$ is closed and convex, where $M_{i}(x)$ is the maximizer set given by
@@
As $\langle g_{i}(x),y_{i}^{*}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}$ is weak-* continuous in $y_{i}^{*}$ for any $x\in\mathtt{X}$ and $\{y_{i}^{*}\in K_{i}^{+};~{}\lVert y_{i}^{*}\rVert_{\mathtt{Y}_{i}^{*}}=1\}$ is weak-* compact due to Alaogu’s theorem, $M_{i}(x)$ is nonempty, and moreover it can be easily verified to be convex and weak-* closed (and hence closed). The convexity of $\{y_{i}^{*}\circ Dg_{i}(x);~{}y_{i}^{*}\in M_{i}(x)\}$ follows from $Dg_{i}(x)\in\mathcal{L}(\mathtt{X},\mathtt{Y}_{i})$ and the convexity of $M_{i}(x)$. To show the closedness, let us assume that $y_{i}^{n^{*}}\circ Dg_{i}(x)\to x^{*}$ in $\mathtt{X}^{*}$ with $y_{i}^{n^{*}}\in M_{i}(x)$. As the latter set is bounded, according to the separability, there exists a subsequence $(y_{i}^{n_{k}^{*}})$ of $(y_{i}^{n^{*}})$ weakly- converging to some $y^{*}\in\mathtt{Y}_{i}^{*}$ (in fact $y^{*}\in M_{i}(x)$ in view that $M_{i}(x)$ is weak- closed). Thus, we can conclude that
@@ 2.3 Constraint Qualifications
Next, we also provide a sufficient condition for the KKT optimality condition to hold. Let $\Pi_{1}(C):=\{\Pi_{1}(x);~{}x\in C\}$. Then, we have the following corollary whose proof is reported in Appendix A:
@@ 3 Extended MFC Problems with Constraints
In this section, we apply the newly obtained FJ optimality condition in Theorem 2.6 to establish the SMP for some C-MFC problems in which we can encompass both dynamic expectations constraints and dynamic pathwise state-control and law constraints. To this end, we first formulate our C-MFC problem and introduce some basic spaces and notations.
Let $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a filtered probability space with the filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\in[0,T]}$ satisfying the usual conditions. For $r,l\in\mathbb{N}$, let $(W_{t})_{t\in[0,T]}$ be a standard $r$-dimensional $(\mathbb{P},\mathbb{F})$-Brownian motion and denote $U\subset\mathbb{R}^{l}$ as the closed and convex control space. We define $\mathcal{X}$ as the space $L^{2}([0,T];\mathbb{R}^{l})/\mathord{\sim}$ equipped with the inner product $\langle\gamma^{1},\gamma^{2}\rangle_{\mathcal{X}}=\int_{0}^{T}\langle\gamma_{t}^{1},\gamma_{t}^{2}\rangle\mathrm{d}t$. Then, $(\mathcal{X},\langle\cdot,\cdot\rangle_{\mathcal{X}})$ is a Hilbert space. For any $\gamma=(\gamma_{t})_{t\in[0,T]}\in\mathcal{X}$, we introduce, for fixed $t\in[0,T]$,
@@
where $\beta^{1}=\beta^{2}$ means that $\beta^{1}(\omega)=\beta^{2}(\omega)$ holds in $\mathcal{X}$ except for a $\mathbb{P}$-null set. It can be easily verified that $\mathscr{U}[0,T]$ is closed and convex in $(\mathbb{L}_{\mathbb{F}}^{2},\langle\cdot,\cdot\rangle_{\mathbb{L}_{\mathbb{F}}^{2}})$. We can observe that $\beta\in L^{2}(\Omega;\mathcal{X})$ for all $\beta\in\mathbb{L}_{\mathbb{F}}^{2}$, and moreover, we have $\|\beta\|_{\mathbb{L}_{\mathbb{F}}^{2}}^{2}:=\mathbb{E}\left[\|\beta\|_{\mathcal{X}}^{2}\right]=\mathbb{E}\left[\langle\beta,\beta\rangle_{\mathcal{X}}\right]$. Let us define $\mathbb{M}^{2}$ ($\subset\mathbb{L}_{\mathbb{F}}^{2}$) as the set of $\mathbb{F}$-adapted (continuous) $\mathbb{R}^{n}$-valued process $\eta=(\eta_{t})_{t\in[0,T]}$ satisfying the Itô’s representation that
@@ 3.1 Partial L-derivative
Let $w:\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})\mapsto\mathbb{R}$. We use $\rho$ to represent a generic element in $\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})$, while $\mu\in\mathcal{P}_{2}(\mathbb{R}^{n})$ and $\nu\in\mathcal{P}_{2}(\mathbb{R}^{l})$ denote its respective marginals. For a probability space $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{P}})$, define the lift $\tilde{w}:L^{2}((\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{P}});\mathbb{R}^{n}\times\mathbb{R}^{l})$ of the element $w$ by $\tilde{w}(\tilde{X},\tilde{\alpha}):=w(\mathcal{L}(\tilde{X},\tilde{\alpha}))$, where $\mathcal{L}(\tilde{X},\tilde{\alpha})$ denotes the joint distribution of $(\tilde{X},\tilde{\alpha})$ under $\tilde{P}$. Then, the L-derivative is defined by
@@
3.2 Formulation of C-MFC problems
In this section, we formulate our C-MFC problems. Let $(b,\sigma):[0,T]\times\mathbb{R}^{n}\times{\cal P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})\times\mathbb{R}^{l}\mapsto\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ be measurable mappings. For any $\alpha=(\alpha_{t})_{t\in[0,T]}\in\mathscr{U}[0,T]$, consider the following controlled Mckean-Vlasov SDE that, $X_{0}^{\alpha}=\kappa\in L^{2}((\Omega,\mathcal{F}_{0},\mathbb{P});\mathbb{R}^{n})$, and for $t\in(0,T]$,
$\mathrm{d}X_{t}^{\alpha}$ $=b(t,X_{t}^{\alpha},\mathcal{L}(X_{t}^{\alpha},\alpha_{t}),\alpha_{t})\mathrm{d}t+\sigma(t,X_{t}^{\alpha},\mathcal{L}(X_{t}^{\alpha},\alpha_{t}),\alpha_{t})\mathrm{d}W_{t}.$ (12)
Let integer $1\leq m_{1}\leq m$. For $\alpha\in\mathscr{U}[0,T]$ and $\phi^{i}:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathcal{X}\mapsto\mathbb{R}$ with $i\in I_{1}:=\{1,2,\ldots,m_{1}\}$, we introduce the dynamic expectation constraints in the sense that, for $i\in I_{1}$,
$\mathbb{E}\left[\phi^{i}\left(t,X_{t}^{\alpha},\mathcal{L}(X_{t}^{\alpha}),\widetilde{\alpha}^{t}\right)\right]\geq 0,\quad\forall t\in[0,T].$ (13)
For $\psi^{j}:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})\times\mathbb{R}^{l}\mapsto\mathbb{R}$ with $j\in I_{2}:=I\setminus I_{1}$, the dynamic pathwise state-control and joint law constraints are formulated by, for $j\in I_{2}$,
$\psi^{j}\left(t,X_{t}^{\alpha},\mathcal{L}(X_{t}^{\alpha},\alpha_{t}),\alpha_{t}\right)\geq 0,\quad\mathrm{d}t\times\mathrm{d}\mathbb{P}\text{-}\mathrm{a.s.}.$ (14)
Here, $X^{\alpha}=(X_{t}^{\alpha})_{t\in[0,T]}$ obeys the controlled Mckean-Vlasov SDE (12) and the process $\widetilde{\alpha}^{t}$ is defined by (7). Then, we introduce the admissible control set $\mathscr{U}_{ad}[0,T]$ as the set of processes $\alpha\in\mathscr{U}[0,T]$ such that there exists $X^{\alpha}=(X_{t}^{\alpha})_{t\in[0,T]}\in\mathbb{M}^{2}$ satisfying (12), (13) and (14). Based on the preparations provided above, we introduce the cost functional as follows, for $\alpha=(\alpha_{t})_{t\in[0,T]}\in\mathscr{U}_{ad}[0,T]\neq\varnothing$,